Random variables z and y described by the PDF if x-+ yo 1 and x.> 0 and y, > 0 0 otherwise a Are x and y independent random variables? b Are they conditionally independent given max(x,y) S 0.5? c Determine the expected value of random variabler, defined byr xy.
Answer #1

The pdf of random variables x,y is

f_{x,y}\left ( x_0,y_0 \right )=\left\{\begin{matrix} K, & x_0+y_0\leqslant 1,x_0,y_0>0\\ 0, & \textup{ otherwise} \end{matrix}\right.

The region is the triangle in the first quadrant between x_0+y_0\leqslant 1,x_0,y_0>0 whose area is \frac{1}{2}\left ( 1\times 1 \right )=0.5 . Hence K=1/0.5=2

The marginal pdfs are

f (zo) 1-10 fr (xo)-Kdyo

Similarly,

f_y\left ( y_0 \right )=K\left ( 1-y_0 \right ),1\leqslant y_0\leqslant 1

a) We can see that f_x\left ( x_0 \right )f_y\left ( y_0 \right)\neq f_{x,y}\left ( x_0,y_0 \right ) , x,y are not independent.

b) The probability P\left ( \max\left ( x_0,y_0 \right ) \leqslant 0.5\right )=\frac{0.5^2}{0.5}=0.5 . Since the region \max\left ( x_0,y_0 \right ) \leqslant 0.5 is square of area 0.25.

P\left (x_0\leqslant 0.5\right )=K\int_{0}^{0.5}\left ( 1-x_0 \right )dx_0\\ P\left (x_0\leqslant 0.5\right )=\frac{3K}{8}=\frac{3}{4}\\

Similarly,

P\left (y_0\leqslant 0.5\right )=\frac{3K}{8}=\frac{3}{4}\\

We can see that  P\left (x_0\leqslant 0.5\right )P\left (y_0\leqslant 0.5\right )\neq P\left ( \max\left ( x_0,y_0 \right ) \leqslant 0.5\right ), x_0,y_0 are not conditionally independent given \max\left ( x_0,y_0 \right ) \leqslant 0.5 .

c) The expected value

E\left ( xy \right )=K\int_{0}^{1}\int_{0}^{1-x_0}x_0y_0dy_0dx_0\\ E\left ( xy \right )=K\int_{0}^{1}x_0\left [ \frac{y_0^2}{2} \right ]_{y_0=0}^{1-x_0}dx_0\\ E\left ( xy \right )=K\int_{0}^{1}x_0\left [ \frac{\left ( 1-x_0 \right )^2}{2} \right ]dx_0\\

E\left ( xy \right )=0.5K\left [\frac{x_0^2}{2} -\frac{2x_0^3}{3}+\frac{x_0^4}{4} \right ]_0^1\\ {\color{Blue}E\left ( xy \right )=\frac{1}{12}}